A local accounting firm employs 24 full-time professionals. The budgeted annual compensation per employee is $45,000. The average chargeable time is 420 hours per client annually. All professional labor costs are included in a single direct-cost category and are allocated to jobs on a per-hour basis.

Other costs are included in a single indirect-cost pool, allocated according to professional labor-hours. Budgeted indirect costs for the year are $790,000, and the firm expects to have 90 clients during the coming year.

What is the budgeted direct labor cost rate per hour? (Round the final answer to the nearest cent.)
Answer #1

Direct labor cost rate per hour =

Total direct labor costs/ total hours

Total direct labor cost:45,000 * 24 = $ 1,080,000

Total hours:420 × 90 = 37,800 hours

Direct labor cost rate per hour = 1,080,000/37,800

= $ 28.5714285714 per hour

=$ 28.57 per hour

The budgeted direct labor cost rate per hour = $ 28.57 per hour.